David likes to party so he decides to head over to Westport in Kansas City for some Halloween shenanigans with his lady friend and some of her friends. At the bar, David’s lady friend orders shots and offers him 2 of the 10 shots she ordered.  5 of the 10 shots have bourbon in them (which David hates). The shots all have a ton of food coloring and sugar in them so it is not possible to visually figure out which ones have bourbon and which ones do not. What is the probability that David is very unlucky and both shots David takes have bourbon in them?
Answer #1

P(Bourbon shot) = 5/10 =1/2
P(2 bourbon shot)=5/10*4/9*1/2=2/9
After the first bourbon shot, there are 9 shots remaining of which 4 are bourbon shots.
Hence the probability that David is unlucky is 2/9.